# GreenScan: Towards large-scale monitoring the health of urban trees using mobile sensing

### 22 Dec 2023 v1

@@ IV-D2 Species-wise tree health analysis
V Limitations and future work
@@
V-1 Feasibility of modelling-based classification
From the correlation matrix in Figure 11, it is seen that is no correlation between CTD and NDVI values. Hence, to develop an autonomous model to classify tree health, both these measured parameters are useful. A scatter plot between NDVI and CTD values for red pine trees in shown in Figure 14. From the scatter plot, it is seen that most of the fair and poor-condition trees are concentrated around a cluster between NDVI (0.20-0.35) and CTD (0-7). Hence, simple white-box machine learning algorithms like SVMs with kernel [54] or logistic regression classifiers [55] can be used to autonomously distinguish between good, poor, and fair condition trees. Further, the methodology can be expanded by adding human-in-the-loop-validation at intermediate steps to enhance the performance of the system.
Figure 14: Scatter plot between NDVI and CTD for red pine trees. The color of the points indicates the ground truth health with red denoting poor, yellow denoting fair, and green denoting good condition trees.
V-2 Direction of movement and Robust positioning system
At present, our methodology is contingent upon aligning the system consisting the imaging sensors with the trees’ orientation. However, given the intended practical application in real-world scenarios, the angle between the tree canopy and the camera’s direction could potentially impact the segmentation of tree canopies. Furthermore, our current positioning method relies on GPS coordinates sourced from the tree survey dataset and the GNSS module within the system, which inherently faces uncertainties in positioning. Thus, an alternative positioning approach that integrates GNSS with computer vision techniques may offer enhanced robustness.
V-3 Real-time segmentation using Mask R-CNN
While the results of the k-fold cross-validation as shown in Table V showcase the reliability of our results, the inference time as shown in Table IV is still high for real-time image processing. Since our system does not need to be instaneous based on the use case, batch processing can be applied after the collection of images at regular intervals in a real-world deployment scenario. For instance, the system can process the captured data while the moving vehicle is waiting at traffic lights.
V-4 Scalability in different weather conditions and geographical boundaries
@@
VI Conclusion
Urban greenery provides various environmental services such as carbon sequestration, and cooling making them essential for building climate-adaptive cities. Currently, urban trees are experiencing atypical amounts of natural and human-induced stresses leading to their volatile health. Yet, high costs make it infeasible for cities to perform frequent inspections on a large scale, leading to adverse health conditions being discovered only after severe damage. The current popular methods for monitoring the health of urban trees rely on an in-person inspection performed by arborists and remote sensing based on satellites or airborne imagery. However, all these methods are riddled with various challenges involving scalability, spatio-temporal resolutions, and quality of assessment. In this work, we developed a novel system to measure tree health autonomously from the ground level in urban cities. The system fuses data from low-cost thermal and multispectral imaging sensors using custom computer vision models optimised for efficiency to generate the tree health indexes namely NDVI and CTD. The custom Mask R-CNN model fine-tuned using transfer learning was employed to fuse the data collected by the imaging sensors on the edge device. Deployment can be performed both in a drive-by sensing paradigm on moving vehicles such as taxis and garbage trucks or in a citizen-science sensing paradigm by humans. Evaluation of the system was performed through data collection experiments in Cambridge, USA. The custom Mask R-CNN developed performed admirably, with an $AP^{IoU=0.50}=0.938$ despite the small dataset used for training. The tree health analysis revealed moderately-strong correlation (r=0.54 with $p-value<0.05$) between our measured NDVI and the remote NDVI obtained from the ground truth dataset. Further, for both the species of trees analysed, our measured NDVI distributions were found to be in theoretical agreement with ground truth tree health conditions. For CTD, a pattern with a theoretical agreement was applicable for one of the species observed. In essence, this work illustrates the potential of autonomous ground-based tree health monitoring on city-wide scales at high temporal resolutions while motivating future research at the intersection of environmental science and computer science.

### 06 Apr 2024 v2

@@ 4.4.2 Species-wise tree health analysis
5 Limitations and future work
@@
5.0.1 Feasibility of modelling-based classification
From the correlation matrix in Figure 11, it is seen that is no correlation between CTD and NDVI values. Hence, in autonomous models to classify tree health, both these measured parameters are useful features. A scatter plot between NDVI and CTD values for red pine trees in shown in Figure 13. From the scatter plot, it is seen that most of the fair and poor-condition trees are concentrated around a cluster between NDVI (0.20-0.35) and CTD (0-7). Hence, simple white-box machine learning algorithms like SVMs (support vector machines) with kernel [63] or logistic regression classifiers [64] can autonomously distinguish between good, and poor or fair condition trees. Further, the methodology can be expanded by adding human-in-the-loop-validation at intermediate steps to enhance the performance of the system.
Figure 13: Scatter plot between NDVI and CTD (in ${}^{\circ}$) for red pine trees. The color of the points indicates the ground truth health with red denoting poor, yellow denoting fair, and green denoting good condition trees.
5.0.2 Direction of movement and Robust positioning system
At present, our methodology is contingent upon aligning the system consisting the imaging sensors with the trees’ orientation. However, given the intended practical application in real-world scenarios, the angle between the tree canopy and the camera’s direction could potentially impact the segmentation of tree canopies. This can be improved by adding a simple image selection algorithm such that the majority of image frame is occupied by pixels belonging to a tree canopy. Furthermore, our current positioning method relies on GPS coordinates sourced from the tree survey dataset and the GNSS module within the system, which inherently faces uncertainties in positioning. Thus, an alternative positioning approach utilising RTK (Real-Time Kinematics) can enhance the positioning robustness.
5.0.3 Scalability in different weather conditions and geographical boundaries
@@
6 Conclusion
Urban greenery provides various environmental services such as carbon sequestration, and cooling making them essential for building climate-adaptive cities. Currently, urban trees are experiencing atypical amounts of natural and human-induced stresses leading to their volatile health. Yet, high costs make it infeasible for cities to perform frequent inspections on a large scale, leading to adverse health conditions being discovered only after severe damage. The current popular methods for monitoring the health of urban trees rely on an in-person inspection performed by arborists and remote sensing based on satellites or airborne imagery. However, all these methods are riddled with various challenges involving scalability, spatio-temporal resolutions, and quality of assessment. In this work, we developed a novel system called GreenScan to measure tree health autonomously from the ground level in urban cities. The GreenScan system fuses data from low-cost thermal and multispectral imaging sensors using custom computer vision models optimised for efficiency to generate the tree health indexes namely NDVI and CTD. The custom Mask R-CNN model fine-tuned using transfer learning was employed to fuse the data collected by the imaging sensors on the edge device. Deployment can be performed both in a drive-by sensing paradigm on moving vehicles such as taxis and garbage trucks or in a citizen-science sensing paradigm by humans. Initial evaluation of the system was performed through data collection experiments in Cambridge, USA. The custom Mask R-CNN developed performed admirably, with an $AP^{IoU=0.50}=0.938$ despite the small dataset used for training. The tree health analysis revealed moderately-strong correlation between our measured NDVI and the remote NDVI obtained from the ground truth dataset. Further, our measured NDVI distributions can be used to flag trees that are specifically not in good health conditions. For the measured CTD, a pattern with a theoretical agreement was applicable for one of the species observed. However, further large-scale evaluation studies over multiple species would help in improving the generalisability of the system. In essence, this work illustrates the potential of autonomous ground-based urban tree health monitoring on city-wide scales at high temporal resolutions and motivates future research at the intersection of environmental science and computer science.